# Exact Gauss-Newton Optimization for Training Deep Neural Networks

### v1

@@ B.1 Generalized Gauss-Newton Hessian Approximation
Show that the generalized Gauss-Newton Hessian approximation scheme for the batch loss (2.3) is $\mathbf{H}^{\mathrm{GN}}=\frac{1}{b}\mathbf{J}^{\top}\mathbf{Q}\mathbf{J}$.
@@
B.2 Gauss-Newton Hessian of the MSE Loss Function
Show that the gradient of the batch MSE loss is $\mathbf{g}=\frac{1}{b}\mathbf{J}^{\top}\mathbf{r}$.
@@ B.2 Gauss-Newton Hessian of the MSE Loss Function
Show that the Hessian of the batch MSE loss is $\mathbf{H}^{\text{GN}}=\frac{1}{b}\mathbf{J}^{\top}\mathbf{J}$.
@@
$\mathbf{H}=\frac{1}{b}\left(\mathbf{J}^{\top}\frac{\partial}{\partial\mathbf{w}}\left[\mathbf{r}\right]+\left(\frac{\partial}{\partial\mathbf{w}}\left[\mathbf{J}^{\top}\right]\right)\mathbf{r}\right)=\frac{1}{b}\left(\mathbf{J}^{\top}\mathbf{J}+\sum_{i=1}^{b}\frac{\partial^{2}}{\partial\mathbf{w}^{2}}\left[\Phi(\mathbf{x}_{i};\mathbf{w})\right]\mathbf{r}_{i}\right),$
@@
B.3 Gauss-Newton Hessian of the Multi-class Cross-entropy Loss Function
Show that the gradient of the batch multi-class cross-entropy loss is $\mathbf{g}=\frac{1}{b}\mathbf{J}^{\top}\mathbf{r}$.
Proof. We recall from Appendix B.1 that the partial derivative of the generic loss function is
@@ B.3 Gauss-Newton Hessian of the Multi-class Cross-entropy Loss Function
where for each sample $i$ in the batch $\mathcal{B}$ we vertically stack Jacobians $\mathbf{J}_{\Phi_{i}}$ as well residuals $\mathbf{r}_{i}$ to form $\mathbf{J}\in{\mathbb{R}}^{bc\times d}$ and $\mathbf{r}\in{\mathbb{R}}^{bc}$. $\Box$
Show that the Hessian of the batch multi-class cross-entropy loss is $\mathbf{H}^{\text{GN}}=\frac{1}{b}\mathbf{J}^{\top}\mathbf{Q}\mathbf{J}$.
Proof. Recall that the GN Hessian for the generalized case (Appendix B.1) is
$\mathbf{H}^{\mathrm{GN}}=\frac{1}{b}\sum_{i=1}^{b}\mathbf{J}_{\Phi_{i}}^{\top}\mathbf{Q}_{\ell_{i}}\mathbf{J}_{\Phi_{i}}.$ (B.1)
@@
B.4 Proof of Lemma 3.2
$\left(b\lambda\mathbf{I}_{d}+\mathbf{J}^{\top}\mathbf{Q}\mathbf{J}\right)^{-1}\mathbf{J}^{\top}\mathbf{Q}=-\left(b\lambda\mathbf{I}_{d}\right)^{-1}\mathbf{J}^{\top}\left(\mathbf{Q}^{-1}+\mathbf{J}\left(b\lambda\mathbf{I}_{d}\right)^{-1}\mathbf{J}^{\top}\right)^{-1}$ (B.2)
Multiplying both sides by $\mathbf{Q}^{-1}\mathbf{r}$ yields
$\left(b\lambda\mathbf{I}_{d}+\mathbf{J}^{\top}\mathbf{Q}\mathbf{J}\right)^{-1}\mathbf{J}^{\top}\mathbf{r}=-\left(b\lambda\mathbf{I}_{d}\right)^{-1}\mathbf{J}^{\top}\left(\mathbf{Q}^{-1}+\mathbf{J}\left(b\lambda\mathbf{I}_{d}\right)^{-1}\mathbf{J}^{\top}\right)^{-1}\mathbf{Q}^{-1}\mathbf{r},$ (B.3)
where the lhs of (B.3) is now equivalent to explicitly solving for $\mathbf{d}^{\mathrm{LM}}$ in Equation (3.1). Simplifying the rhs of (B.3) results in
$\mathbf{d}^{\mathrm{LM}}=-\mathbf{J}^{\top}\left(b\lambda\mathbf{Q}^{-1}+\mathbf{J}\mathbf{J}^{\top}\right)^{-1}\mathbf{Q}^{-1}\mathbf{r}.$ (B.4)
$\left(b\lambda\mathbf{Q}^{-1}+\mathbf{J}\mathbf{J}^{\top}\right)^{-1}\mathbf{Q}^{-1}=\left(\mathbf{Q}\left(b\lambda\mathbf{Q}^{-1}+\mathbf{J}\mathbf{J}^{\top}\right)\right)^{-1}=\left(b\lambda\mathbf{I}_{bc}+\mathbf{Q}\mathbf{\mathbf{J}}\mathbf{J}^{\top}\right)^{-1}.$ (B.5)
So that the direction can be calculated as
$\mathbf{d}^{\mathrm{LM}}=-\mathbf{J}^{\top}\left(\mathbf{Q}\mathbf{J}\mathbf{J}^{\top}+b\lambda\mathbf{I}_{bc}\right)^{-1}\mathbf{r},$ (B.6)
which corresponds to the procedure outlined in Algorithm 1. $\Box$
B.5 Proof of Lemma 4.1
Proof. From Assumption 4.1, we have
$={\mathcal{L}}_{N}(\mathbf{w}_{t})-\alpha_{t}\left\langle\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t}),\left(\frac{1}{b}\mathbf{J}_{t}^{\top}\mathbf{Q}_{t}\mathbf{J}_{t}+\lambda_{t}\mathbf{I}\right)^{-1}\mathbf{g}_{t}\right\rangle+\frac{\alpha_{t}^{2}L_{1}}{2}\left\lVert\left(\frac{1}{b}\mathbf{J}_{t}^{\top}\mathbf{Q}_{t}\mathbf{J}_{t}+\lambda_{t}\mathbf{I}\right)^{-1}\mathbf{g}_{t}\right\rVert^{2}.$ (B.7)
Next, set $\mathbf{B}_{t}=\left(\frac{1}{b}\mathbf{J}_{t}^{\top}\mathbf{Q}_{t}\mathbf{J}_{t}+\lambda_{t}\mathbf{I}\right)^{-1}$ in Assumption 4.4. From Assumption 4.3, we can find $\mathbf{B}_{t}\succeq\mu\mathbf{I}$ with $\mu\coloneqq b/(b\lambda_{t}+\bar{\kappa}\bar{\sigma}^{2})$. Now, using these and Assumption 4.3 in (B.7), we have
${\mathcal{L}}_{N}(\mathbf{w}_{t+1})$ $\leq{\mathcal{L}}_{N}(\mathbf{w}_{t})-\alpha_{t}\left\langle\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t}),\mathbf{B}_{t}\mathbf{g}_{t}\right\rangle+\frac{b^{2}\alpha_{t}^{2}L_{1}}{2(b\lambda_{t}+\underaccent{\bar}{\kappa}\underaccent{\bar}{\sigma}^{2})^{2}}\|\mathbf{g}_{t}\|^{2}.$ (B.8)
Notice that by Assumption 4.2, we get that ${\mathbb{E}}[\mathbf{g}_{t}|\mathbf{w}_{t}]=\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})$, and hence the inequality in Assumption 4.4 may be written as ${\mathbb{E}}\left[\langle\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t}),\mathbf{B}_{t}\mathbf{g}_{t}\rangle|\mathbf{w}_{t}\right]\geq\mu K\|\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}$. Taking conditional expectation on both sides of (B.8) with respect to $\xi$ and using Assumption 4.4, we get
${\mathbb{E}}[{\mathcal{L}}_{N}(\mathbf{w}_{t+1})|\mathbf{w}_{t}]\leq{\mathcal{L}}_{N}(\mathbf{w}_{t})-\frac{3\alpha_{t}}{5\lambda_{t}}\|\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}+\frac{b^{2}\alpha_{t}^{2}L_{1}}{2(b\lambda_{t}+\underaccent{\bar}{\kappa}\underaccent{\bar}{\sigma}^{2})^{2}}{\mathbb{E}}[\|\mathbf{g}_{t}\|^{2}|\mathbf{w}_{t}].$ (B.9)
We have also, using Assumption 4.2,
$={\mathbb{E}}[\|\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}|\mathbf{w}_{t}]+2{\mathbb{E}}[\langle\mathbf{g}_{t}-\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t}),\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\rangle|\mathbf{w}_{t}]$
$\qquad+{\mathbb{E}}[\|\mathbf{g}_{t}-\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}|\mathbf{w}_{t}]$
$\leq\|\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}+\frac{\sigma_{g}^{2}}{b}.$ (B.10)
Now, using (B.10) in (B.9), we get
${\mathbb{E}}[{\mathcal{L}}_{N}(\mathbf{w}_{t+1})|\mathbf{w}_{t}]$ $\leq{\mathcal{L}}_{N}(\mathbf{w}_{t})-\left(\frac{3\alpha_{t}}{5\lambda_{t}}-\frac{b^{2}\alpha_{t}^{2}L_{1}}{2(b\lambda_{t}+\underaccent{\bar}{\kappa}\underaccent{\bar}{\sigma}^{2})^{2}}\right)\|\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}+\frac{b\sigma_{g}^{2}\alpha_{t}^{2}L_{1}}{2(b\lambda_{t}+\underaccent{\bar}{\kappa}\underaccent{\bar}{\sigma}^{2})^{2}}.$ (B.11)
Using $\alpha_{t}\leq(b\lambda_{t}+\underaccent{\bar}{\kappa}\underaccent{\bar}{\sigma}^{2})^{2}/(b^{2}\lambda_{t}L_{1})$ in (B.11) gives (4.5). $\Box$
B.6 Proof of Theorem 4.2
Proof. Taking conditional expectation on both sides of (4.5), we have
${\mathbb{E}}[\|\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}]\leq\frac{10\lambda}{\alpha_{t}}\left({\mathbb{E}}[{\mathcal{L}}_{N}(\mathbf{w}_{t})]-{\mathbb{E}}[{\mathcal{L}}_{N}(\mathbf{w}_{t+1})]\right)+\frac{5\sigma_{g}^{2}}{b}.$ (B.12)
Summing (B.12) over $t=0,1,\ldots,T-1$, and using Assumption 4.1, we get
$\sum_{t=0}^{T-1}{\mathbb{E}}[\|\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}]$ $\leq 10\lambda\sum_{t=0}^{T-1}\frac{1}{\alpha_{t}}\left({\mathbb{E}}[{\mathcal{L}}_{N}(\mathbf{w}_{t})]-{\mathbb{E}}[{\mathcal{L}}_{N}(\mathbf{w}_{t+1})]\right)+\frac{5\sigma_{g}^{2}T}{b}$
$=\frac{10\lambda}{\alpha_{0}}{\mathcal{L}}_{N}(\mathbf{w}_{0})+10\lambda\sum_{t=1}^{T-1}\left(\frac{1}{\alpha_{t}}-\frac{1}{\alpha_{t-1}}\right){\mathbb{E}}[{\mathcal{L}}_{N}(\mathbf{w}_{t})]-\frac{10\lambda}{\alpha_{T-1}}{\mathbb{E}}[{\mathcal{L}}_{N}(\mathbf{w}_{T+1})]$
$\qquad+\frac{5\sigma_{g}^{2}T}{b}$
$\leq\frac{10\lambda}{\alpha_{0}}\bar{C}+10\lambda\bar{C}\sum_{t=1}^{T-1}\left(\frac{1}{\alpha_{t}}-\frac{1}{\alpha_{t-1}}\right)-\frac{10\lambda}{\alpha_{T-1}}{\mathcal{L}}_{N}^{*}+\frac{5\sigma_{g}^{2}T}{b}$
$=\frac{10\lambda}{\alpha_{T-1}}(\bar{C}-{\mathcal{L}}_{N}^{*})+\frac{5\sigma_{g}^{2}T}{b}$
$\leq\frac{4b^{2}\lambda^{2}L_{1}}{\bar{\kappa}\bar{\sigma}^{2}(b\lambda+\underaccent{\bar}{\kappa}\underaccent{\bar}{\sigma}^{2})}(\bar{C}-{\mathcal{L}}_{N}^{*})+\frac{5\sigma_{g}^{2}T}{b},$ (B.13)
where we have used that $\alpha_{T-1}\geq(5\bar{\kappa}\bar{\sigma}^{2}(b\lambda+\underaccent{\bar}{\kappa}\underaccent{\bar}{\sigma}^{2}))/(2b^{2}\lambda L_{1})$ holds under the condition $b\geq\lambda^{-1}((5/2)\bar{\kappa}\bar{\sigma}-\underaccent{\bar}{\kappa}\underaccent{\bar}{\sigma})>0$. Dividing (B.13) through by $T$ gives the result. $\Box$
Appendix C Algorithms
Algorithm 2 Calculate direction using x2 QR factorization (MSE loss)
@@ Appendix C Algorithms
Momentum
One aspect of second-order methods not yet addressed is issue (d), specifically, the noise in Hessian estimates. Consider the direction $\mathbf{d}_{t}=-\mathbf{H}^{-1}_{t}\mathbf{g}_{t}$, where both the (approximate) Hessian $\mathbf{H}_{t}$ and the gradient $\mathbf{g}_{t}$ are stochastic estimates of the true derivatives of the empirical risk (2.2). By conditioning the gradient with a noisy Hessian inverse we can introduce additional error in the descent direction. A common technique to stabilize the updates and speed up convergence is temporal averaging (or momentum) that consists in combining past and current estimates. For ML applications the common variants include simple accumulation [22], exponential moving average (EMA) [30], bias-corrected EMA [36, 78] and momentum with an extrapolation step [51].
For diagonal scaling methods, including first-order accelerated algorithms [36, 22, 30] and SOMs [78, 42], the accumulated estimates of both first and second moments are kept separately resulting in $\mathcal{O}\left(d\right)$ space complexity. For Gauss-Newton methods we could alternatively reduce the variance of the Jacobian $\mathbf{J}$, e.g., with SVRG [35], SAGA [19] or SARAH [52], which results in a space complexity of $\mathcal{O}\left(bcd\right)$. Another option is to apply momentum to the descent direction $\mathbf{d}_{t}$, explored in [72, 37], which requires storing a vector of size $d$. Since we never explicitly materialize neither the Hessian nor the gradient (Algorithm 1), we follow the latter approach with bias-corrected EMA by default.
Line Search
Instead of setting $\alpha$ and $\lambda$ manually, we can make these hyper-parameters adapt to the training step to eliminate the costly procedure of hyper-parameter tuning. Line search is a popular technique in deterministic optimization [53] which consists in iteratively searching for a learning rate satisfying some minimum criteria (e.g., Wolfe’s conditions). Providing the same theoretical guarantees becomes challenging in the stochastic setting due to noisy gradients, as outlined in [17]. Despite these difficulties, line search has been successfully applied in practice to both stochastic FOMs [72, 57] and SOMs [59, 75]. In EGN we adopt the strategy suggested by [72], which allows to minimize the time spent in evaluating the loss by incorporating a reset strategy in the beginning of each search (Algorithm 3).
@@
Adaptive regularization
Adaptive regularization techniques for stochastic SOMs are explored in [46, 37, 60, 59] modifying the original Levenberg-Marquardt rule to the stochastic setting. The idea is to track $\rho$, defined as the ratio between the decrease in the actual loss function (2.3) and the decrease in the quadratic model $\mathcal{M}\left(\Delta\mathbf{w}\right)=\mathcal{L}_{b}\left(\mathbf{y}_{t},\Phi(\mathbf{x}_{t};\mathbf{w}_{t})\right)+\mathbf{g}_{t}^{\top}\Delta\mathbf{w}+\frac{1}{2b}\Delta\mathbf{w}^{\top}\mathbf{J}_{t}^{\top}\mathbf{Q}_{t}\mathbf{J}_{t}\Delta\mathbf{w}$, where $\Delta\mathbf{w}=\mathbf{w}-\mathbf{w}_{t}$. This yields
$\rho:=\frac{\mathcal{L}_{b}\left(\mathbf{y}_{t},\Phi(\mathbf{x}_{t};\mathbf{w}_{t+1})\right)-\mathcal{L}_{b}\left(\mathbf{y}_{t},\Phi(\mathbf{x}_{t};\mathbf{w}_{t})\right)}{\mathbf{g}_{t}^{\top}\Delta\mathbf{w}+\frac{1}{2b}\Delta\mathbf{w}^{\top}\mathbf{J}_{t}^{\top}\mathbf{Q}_{t}\mathbf{J}_{t}\Delta\mathbf{w}},$ (C.1)
which measures the accuracy of the quadratic model. In case $\rho$ is small or negative, $\mathcal{M}\left(\Delta\mathbf{w}\right)$ provides inaccurate approximation and the value $\lambda$ is increased. In case $\rho$ is large, $\lambda$ is decreased to give more weight to $\mathcal{M}\left(\Delta\mathbf{w}\right)$ (Algorithm 4). As empirically found by [37], compared to the deterministic LM the increase/decrease coefficients need to be less aggressive to reduce the oscillations of $\lambda_{t}$.
@@
Algorithm 5 EGN
2:  Initialize momentum: $\mathbf{m}_{0}=0$
[…]
4:     Sample a mini-batch $\mathcal{B}_{t}$ from $\mathcal{D}$
5:     Estimate $\mathbf{r}_{t}$ and $\mathbf{J}_{t}$ (e.g., via backpropagation)
6:     Find direction $\mathbf{d}_{t}$ via Algorithm 1
9:     Line search for $\alpha_{t}$ via Algorithm 3
11:     Update $\lambda_{t+1}$ via Algorithm 4
12:  end for
13:  Return $\mathbf{w}_{t}$
Appendix D Experiment Details
D.1 Supervised Learning
IMDB Reviews [45] is a TFDS [68] dataset that contains $25000$ train samples and $25000$ test samples of movie reviews in a text format. Before passing the samples to the model $\Phi$, we pre-process the raw text data with spaCy [32] en_core_web_lg pipeline that converts a text review into a $300$-dimensional vector of numbers.
The optimal sets of hyper-parameters are presented in Table 4.
@@ D.1 Supervised Learning
D.2 Learning LQR Controllers
@@
4:     Given $\mathbf{K}_{p-1}$ estimate the corresponding weights $\mathbf{w}$ through a policy evaluation algorithm, eg. Algorithm 7
@@
8:     Convert $[S_{i},A_{i}]$ and $[S_{i+1},A^{\prime}]$ to quadratic feature vectors $\mathbf{x}$ and $\mathbf{x^{\prime}}$
@@
System matrices for the model of a binary distillation tower (BDT) [18] are defined as follows:
$\tiny\mathbf{A}=\left[\begin{array}[]{ccccccccccc}-0.014&0.0043&0&0&0&0&0&0&0&0&0\\ 0.0095&-0.0138&0.0046&0&0&0&0&0&0&0&0.0005\\ 0&0.0095&-0.0141&0.0063&0&0&0&0&0&0&0.0002\\ 0&0&0.0095&-0.0158&0.011&0&0&0&0&0&0\\ 0&0&0&0.0095&-0.0312&0.015&0&0&0&0&0\\ 0&0&0&0&0.0202&-0.0352&0.022&0&0&0&0\\ 0&0&0&0&0&0.0202&-0.0422&0.028&0&0&0\\ 0&0&0&0&0&0&0.0202&-0.0482&0.037&0&0.0002\\ 0&0&0&0&0&0&0&0.0202&-0.0572&0.042&0.0005\\ 0&0&0&0&0&0&0&0&0.0202&-0.0483&0.0005\\ 0.0255&0&0&0&0&0&0&0&0&0.0255&-0.0185\end{array}\right]$
$\tiny\mathbf{B}=\left[\begin{array}[]{ccc}0&0&0\\ 0.000005&-0.00004&0.0025\\ 0.000002&-0.00002&0.005\\ 0.000001&-0.00001&0.005\\ 0&0&0.005\\ 0&0&0.005\\ -0.000005&0.00001&0.005\\ -0.00001&0.00003&0.005\\ -0.00004&0.000005&0.0025\\ -0.00002&0.000002&0.0025\\ 0.00046&0.00046&0\end{array}\right].$
The matrices represent a continuous system. We discretize the system using a Zero-Order Hold (ZOH) method with a sampling rate of $\Delta T=0.1s$.
@@
System matrices for the linearized vertical plane dynamics of an aircraft (UAV) [34] are defined as follows:
$\mathbf{A}=\left[\begin{array}[]{ccccc}0&0&1.132&0&-1\\ 0&-0.0538&-0.1712&0&0.0705\\ 0&0&0&1&0\\ 0&0.0485&0&-0.8556&-1.013\\ 0&-0.2909&0&1.0532&-0.6859\end{array}\right]\quad\mathbf{B}=\left[\begin{array}[]{ccc}0&0&0\\ -0.12&1&0\\ 0&0&0\\ 4.419&0&-1.665\\ 1.575&0&-0.0732\end{array}\right].$
The matrices represent a continuous system. We discretize the system using a Zero-Order Hold (zoh) method with a sampling rate of $\Delta T=0.1s$.
@@
Table 5: Optimal Hyper-parameters for BDT and UAV
BDT
SGD 0.0000005 - - - -
Adam 0.1 - - - -
EGN 1.0 1.0 0.0 False -
SGN 1.0 1.0 - - 10
UAV
SGD 0.00008 - - - -
Adam 0.02 - - - -
EGN 0.2 1.0 0.0 False -
SGN 1.0 1.0 - - 10
D.3 Reinforcement Learning with DQN

### v2

@@ A.1 Generalized Gauss-Newton Hessian Approximation
The generalized Gauss-Newton Hessian approximation scheme for the batch loss (3.3) is $\mathbf{H}^{\mathrm{GN}}=\frac{1}{b}\mathbf{J}^{\top}\mathbf{Q}\mathbf{J}$.
@@
A.2 Gauss-Newton Hessian of the MSE Loss Function
Claim
The gradient of the batch MSE loss is $\mathbf{g}=\frac{1}{b}\mathbf{J}^{\top}\mathbf{r}$.
@@ A.2 Gauss-Newton Hessian of the MSE Loss Function
Claim
The Hessian of the batch MSE loss is $\mathbf{H}^{\text{GN}}=\frac{1}{b}\mathbf{J}^{\top}\mathbf{J}$.
@@
$\mathbf{H}=\frac{1}{b}\left(\mathbf{J}^{\top}\frac{\partial}{\partial\mathbf{w}}\left[\mathbf{r}\right]+\left(\frac{\partial}{\partial\mathbf{w}}\left[\mathbf{J}^{\top}\right]\right)\mathbf{r}\right)=\frac{1}{b}\left(\mathbf{J}^{\top}\mathbf{J}+\sum_{i=1}^{b}\frac{\partial^{2}}{\partial\mathbf{w}^{2}}\left[\Phi(\mathbf{x}_{i};\mathbf{w})\right]\mathbf{r}_{i}\right).$
@@
A.3 Gauss-Newton Hessian of the Multi-class Cross-entropy Loss Function
Claim
The gradient of the batch multi-class cross-entropy loss is $\mathbf{g}=\frac{1}{b}\mathbf{J}^{\top}\mathbf{r}$.
Proof. We recall from A.1 that the partial derivative of the generic loss function is
@@ A.3 Gauss-Newton Hessian of the Multi-class Cross-entropy Loss Function
where for each sample $i$ in the batch $\mathcal{B}$ we vertically stack Jacobians $\mathbf{J}_{\Phi_{i}}$ as well as residuals $\mathbf{r}_{i}$ to form $\mathbf{J}\in{\mathbb{R}}^{bc\times d}$ and $\mathbf{r}\in{\mathbb{R}}^{bc}$. $\Box$
Claim
The Hessian of the batch multi-class cross-entropy loss is $\mathbf{H}^{\text{GN}}=\frac{1}{b}\mathbf{J}^{\top}\mathbf{Q}\mathbf{J}$.
Proof. Recall that the GN Hessian for the generalized case (A.1) is
$\mathbf{H}^{\mathrm{GN}}=\frac{1}{b}\sum_{i=1}^{b}\mathbf{J}_{\Phi_{i}}^{\top}\mathbf{Q}_{\ell_{i}}\mathbf{J}_{\Phi_{i}}.$ (A.1)
@@
Appendix B Algorithms
Algorithm 3 Calculate direction using QR factorization (MSE loss)
@@ Appendix B Algorithms
Appendix C Experiment Details
C.1 Supervised Learning
@@
SQN 0.02 - 0.0 - -
IMDB Reviews [52] is a TFDS [75] dataset that contains $25000$ training samples and $25000$ testing samples of movie reviews in a text format. Before passing the samples to the model $\Phi$, we pre-process the raw text data with spaCy [38] en_core_web_lg pipeline which converts a text review into a $300$-dimensional vector of numbers.
The optimal sets of hyper-parameters are presented in Table 5.
C.2 Learning LQR Controllers
Table 6: Optimal Hyper-parameters for BDT and UAV
BDT
SGD 0.0000005 - - - -
Adam 0.1 - - - -
EGN 1.0 1.0 0.0 False -
SGN 1.0 1.0 - - 10
UAV
SGD 0.00008 - - - -
Adam 0.02 - - - -
EGN 0.2 1.0 0.0 False -
SGN 1.0 1.0 - - 10
Table 7: Optimal Hyper-parameters for Acrobot-v1 and Freeway-v1
Acrobot-v1
SGD 0.001 - - - -
Adam 0.0003 - - - -
EGN 0.1 1.0 0.0 False -
SGN 0.005 1.0 - - 3
Freeway-v1
SGD 0.1 - - - -
Adam 0.0003 - - - -
EGN 0.4 1.0 0.0 False -
SGN 0.5 1.0 - - 5
@@ C.2 Learning LQR Controllers
4:  Given $\mathbf{K}_{p-1}$ estimate the corresponding weights $\mathbf{w}$ through a policy evaluation algorithm, e.g., Algorithm 7
@@
8:  Convert $S_{i},A_{i}]$ and $S_{i+1},A^{\prime}]$ to quadratic feature vectors $\mathbf{x}$ and $\mathbf{x^{\prime}}$
@@
System matrices for the model of a binary distillation tower (BDT) follow [19]. The matrices represent a continuous system. We discretize the system using a Zero-Order Hold (ZOH) method with a sampling rate of $\Delta T=0.1s$.
@@
System matrices for the linearized vertical plane dynamics of an aircraft (UAV) are taken from [40]. The matrices represent a continuous system. We discretize the system using a Zero-Order Hold (zoh) method with a sampling rate of $\Delta T=0.1s$.
@@
C.3 Reinforcement Learning with DQN